# Shaocheng Liu and Qi Chen

### v2

@@ IV-F Entropy functions on the faces involving $\hat{U}_{3,5}^{1}$
For $F=(\hat{U}_{3,5}^{1},U_{1,3}^{234})$ , $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a=\log k$ for integer $k>0$.
@@
$H(X_{i\cup K})+H(X_{j\cup K})$ $=H(X_{K})+H(X_{ij\cup K}),\mid K\mid=2,\{1\}\subseteq K\subseteq N_{4}.$
For $(x_{i},i\in N_{4})\in\mathcal{X}_{N_{4}}$ with $p(x_{1234})>0$, above information equalties imply that the probability mass function satisfies
@@
$n^{(1)}=n^{(2)}=\dots=n^{(t)}$ (IV.F.36)
and so they can be simplified to $k$. Then by (IV.F.32),
$a$ $=\sum_{j=1}^{t}p_{j}\log k=\log k\sum_{j=1}^{t}p_{j}=\log k.$ (IV.F.37)
Hence $a$ can just take the value of $\log k$, where $k$ is a positive integer.
As for “if” part of the theorem, it can be proved by Lemma 2 and the fact that $a=\log{k}$ on the ray $\hat{U}^{1}_{3,5}$ and whole ray $U_{1,3}^{234}$ are entropic. ∎
@@
For $F=(\hat{U}_{3,5}^{1},U_{2,3}^{123})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a=\log{k_{1}},b=\log{k_{2}}$ for positive integer $k_{1},k_{2}$.
@@
$H(X_{ik})+H(X_{jk})$ $=H(X_{k})+H(X_{i,j,k}),\{i,j,k\}=\{2,3,4\},$
$H(X_{i\cup K})+H(X_{j\cup K})$ $=H(X_{K})+H(X_{ij\cup K}),K=\{1,2\},\{1,3\}.$
For $(x_{i},i\in N_{4})\in\mathcal{X}_{N_{4}}$ with $p(x_{1234})>0$, above information equalties imply that the probability mass function satisfies
$p(x_{1},x_{2},x_{3},x_{4})$ $=p(x_{1},x_{2},x_{3})$ (IV.F.38)
$=p(x_{1},x_{2},x_{4})$ (IV.F.39)
$=p(x_{1},x_{3},x_{4})$ (IV.F.40)
$=p(x_{2},x_{3},x_{4}),$ (IV.F.41)
$p(x_{1},x_{2})$ $=p(x_{1})p(x_{2}),$ (IV.F.42)
$p(x_{1},x_{3})$ $=p(x_{1})p(x_{3}),$ (IV.F.43)
$p(x_{1},x_{4})$ $=p(x_{1})p(x_{4}),$ (IV.F.44)
$p(x_{2},x_{3})$ $=p(x_{2})p(x_{3}),$ (IV.F.45)
$p(x_{2},x_{4})$ $=p(x_{2})p(x_{4}),$ (IV.F.46)
$p(x_{3},x_{4})$ $=p(x_{3})p(x_{4}),$ (IV.F.47)
$p(x_{2},x_{3})p(x_{2},x_{4})$ $=p(x_{2})p(x_{2},x_{3},x_{4}),$ (IV.F.48)
$p(x_{2},x_{3})p(x_{3},x_{4})$ $=p(x_{3})p(x_{2},x_{3},x_{4}),$ (IV.F.49)
$p(x_{2},x_{4})p(x_{3},x_{4})$ $=p(x_{4})p(x_{2},x_{3},x_{4}),$ (IV.F.50)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{2},x_{4})$ $=p(x_{1},x_{2})p(x_{1},x_{2},x_{3},x_{4}),$ (IV.F.51)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{3},x_{4})$ $=p(x_{1},x_{3})p(x_{1},x_{2},x_{3},x_{4}).$ (IV.F.52)
By (IV.F.38)-(IV.F.41), (IV.F.51) and (IV.F.52), we have
$p(x_{1},x_{2},x_{3},x_{4})=p(x_{1},x_{2})=p(x_{1},x_{3}).$ (IV.F.53)
@@
$p(x_{1},x_{2},x_{3},x_{4})=p(x_{1})p(x_{2})=p(x_{1})p(x_{3}),$ (IV.F.54)
which implies $p(x_{2})=p(x_{3})$. Since $X_{2}$ and $X_{3}$ are independent by (IV.F.45), with Lemma 1, $X_{2}$ and $X_{3}$ are uniformly distributed on $\mathcal{X}_{2}$ and $\mathcal{X}_{3}$ and
$H(X_{2})=H(X_{3})=\log k,$ (IV.F.55)
where $k=|\mathcal{X}_{2}|=|\mathcal{X}_{3}|$. Note that $X_{2}$ is independent of $X_{3}$ and $X_{4}$, replacing $p(x_{2},x_{3})$ and $p(x_{2},x_{4})$ by $p(x_{2})p(x_{3})$ and $p(x_{2})p(x_{4})$ in (IV.F.48), respectively, we have
$p(x_{2},x_{3},x_{4})=p(x_{2})p(x_{3})p(x_{4}).$ (IV.F.56)
Since $p(x_{1},x_{2},x_{3},x_{4})=p(x_{2},x_{3},x_{4})$ by (IV.F.41), together with (IV.F.54) and (IV.F.56),
$p(x_{1})=p(x_{3})p(x_{4})=p(x_{2})p(x_{4}).$ (IV.F.57)
Note that $p(x_{2})=p(x_{3})=\frac{1}{k}$, we obtain
$p(x_{1})=\dfrac{1}{k}p(x_{4}).$ (IV.F.58)
@@
$p(x_{1})=\dfrac{1}{k}p(x_{4}^{\prime}).$ (IV.F.60)
Equating (IV.F.58) and (IV.F.60),
$p(x_{4})=p(x_{4}^{\prime}),$ (IV.F.61)
which implies that $X_{4}$ is uniformly distributed on $\mathcal{X}_{4}$. Similarly, $X_{1}$ is uniformly distributed on $\mathcal{X}_{1}$. Assume $|X_{4}|=k_{1}$, we conclude that
$H(X_{1})=\log kk_{1},H(X_{4})=\log k_{1}.$ (IV.F.62)
@@
$p(x_{1},x_{4})=p(x_{1})p(x_{4})=\dfrac{1}{kk_{1}}\times\dfrac{1}{k_{1}}=\dfrac{1}{kk^{2}_{1}}.$ (IV.F.63)
@@
$p(x_{1},x_{4})$ $=\sum_{x_{2},x_{3}}p(x_{1},x_{2},x_{3},x_{4})=\sum_{x_{2},x_{3}}p(x_{2},x_{3},x_{4})$ (IV.F.64)
$=\sum_{x_{2},x_{3}}p(x_{2})p(x_{3})p(x_{4})=\dfrac{k_{2}}{k^{2}k_{1}},$ (IV.F.65)
where $k_{2}=\mid\{(x_{2},x_{3})|\quad p(x_{1},x_{2},x_{3},x_{4})>0\}\mid$, we conclude that $k=k_{1}k_{2}$. As $\mathbf{h}\in F$ and $(X_{i},i\in N_{4})$ is its characterizing random vector, we have
$H(X_{1})$ $=2a+b,$ (IV.F.66)
$H(X_{2})$ $=H(X_{3})=a+b,$ (IV.F.67)
$H(X_{4})$ $=a.$ (IV.F.68)
Together with (IV.F.55) and (IV.F.62), we obtain
$a=\log k_{1},b=\log\frac{k}{k_{1}}=\log k_{2},$ (IV.F.69)
which forms an outer bound on the entropy region on $F$, that is, $a$ and $b$ must take the value of $\log k_{1}$ and $\log k_{2}$, respectively, where $k_{1}$ and $k_{2}$ are positive integers.
The ”if” part of the theorem can be immediately proved using Lemma 2, given the fact that $a=\log{k_{1}}$ in $\hat{U}^{1}_{3,5}$, and $b=\log{k_{2}}$ in $U_{2,3}^{123}$, both of which are entropic. ∎
@@
For $F=(\hat{U}_{3,5}^{1},\mathcal{W}_{2}^{23})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a+b$= $\log{k}$ for integer $k>0$.
@@
$H(X_{i\cup K})+H(X_{j\cup K})$ $=H(X_{K})+H(X_{ij\cup K}),\mid K\mid=2,\{1\}\subseteq K.$
For $(x_{i},i\in N_{4})\in\mathcal{X}_{N_{4}}$ with $p(x_{1234})>0$, above information equalties imply that the probability mass function satisfies
$p(x_{1},x_{2},x_{3},x_{4})$ $=p(x_{1},x_{2},x_{3})$ (IV.F.70)
$=p(x_{1},x_{2},x_{4})$ (IV.F.71)
$=p(x_{1},x_{3},x_{4})$ (IV.F.72)
$=p(x_{2},x_{3},x_{4}),$ (IV.F.73)
$p(x_{1},x_{2})$ $=p(x_{1})p(x_{2}),$ (IV.F.74)
$p(x_{1},x_{3})$ $=p(x_{1})p(x_{3}),$ (IV.F.75)
$p(x_{1},x_{4})$ $=p(x_{1})p(x_{4}),$ (IV.F.76)
$p(x_{2},x_{4})$ $=p(x_{2})p(x_{4}),$ (IV.F.77)
$p(x_{3},x_{4})$ $=p(x_{3})p(x_{4}),$ (IV.F.78)
$p(x_{2},x_{3})p(x_{2},x_{4})$ $=p(x_{2})p(x_{2},x_{3},x_{4}),$ (IV.F.79)
$p(x_{2},x_{3})p(x_{3},x_{4})$ $=p(x_{3})p(x_{2},x_{3},x_{4}),$ (IV.F.80)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{2},x_{4})$ $=p(x_{1},x_{2})p(x_{1},x_{2},x_{3},x_{4}),$ (IV.F.81)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{3},x_{4})$ $=p(x_{1},x_{3})p(x_{1},x_{2},x_{3},x_{4}),$ (IV.F.82)
$p(x_{1},x_{2},x_{4})p(x_{1},x_{3},x_{4})$ $=p(x_{1},x_{4})p(x_{1},x_{2},x_{3},x_{4}).$ (IV.F.83)
By (IV.F.70)-(IV.F.73) and (IV.F.81)-(IV.F.83), we have
$p(x_{1},x_{2},x_{3},x_{4})=p(x_{1},x_{2})=p(x_{1},x_{3})=p(x_{1},x_{4}).$ (IV.F.84)
Since $X_{1}$ is independent of $X_{2},X_{3}$ and $X_{4}$ by (IV.F.74)-(IV.F.76), together with(IV.F.84)
$p(x_{1})p(x_{2})=p(x_{1})p(x_{3})=p(x_{1})p(x_{4}),$ (IV.F.85)
which implies $p(x_{2})=p(x_{3})=p(x_{4})$. Note that $X_{4}$ is independent of $X_{2}$ and $X_{3}$ by (IV.F.77) and (IV.F.78), by Lemma 1, $X_{i}$ is uniformly distributed on $\mathcal{X}_{i}$ and $H(X_{i})=\log k$ for $i=2,3,4$, where $k=|\mathcal{X}_{2}|=|\mathcal{X}_{3}|=|\mathcal{X}_{4}|$. As $\mathbf{h}\in F$ and $(X_{i},i\in N_{4})$ is its characterizing random vector, we have
$H(X_{2})=H(X_{3})=H(X_{4})=a+b.$ (IV.F.86)
Hence $a+b$ must take the value of $\log k$.
As for the “if” part of the theorem, we show that all the polymatroids $(a,b)$ satisfying $a+b=\log k$ are entropic. Let $p_{1},p_{2},\dots,p_{k}$ be positive value satisfying $H(p_{1},p_{2},...p_{k})=a$. Let $X_{1}$ be distributed on $\{0,1,2,..k^{2}-1\}$ such that
$p_{X_{1}}(i)=\dfrac{p_{\lceil\frac{i+1}{k}\rceil}}{k},\text{ for}i=0,1,...,k^{2}-1.$ (IV.F.87)
Let $X_{2}$ be independent of $X_{1}$ and uniformly distributed on $\{0,1,...,k-1\}$. Let
$X_{3}$ $=(\lfloor\frac{X_{1}}{k}\rfloor+X_{2})\mod k,$ (IV.F.88)
$X_{4}$ $=(X_{1}\mod k+X_{2})\mod k.$ (IV.F.89)

### v3

@@ IV-F Entropy functions on the faces involving $\hat{U}_{3,5}^{1}$
For $F=(\hat{U}_{3,5}^{1},U_{1,3}^{234})$ , $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a=\log v$ for integer $v>0$.
@@
$H(X_{i\cup K})+H(X_{j\cup K})$ $=H(X_{K})+H(X_{ij\cup K}),\mid K\mid=2,$
$\qquad\qquad\qquad\qquad\{1\}\subseteq K\subseteq N_{4}.$
For $(x_{i},i\in N_{4})\in\mathcal{X}_{N_{4}}$ with $p(x_{1234})>0$, above information equalities imply that the probability mass function satisfies
@@
$n^{(1)}=n^{(2)}=\dots=n^{(t)}$ (IV.F.35)
and so they can be simplified to $v$. Then by (IV.F.31),
$a$ $=\sum_{j=1}^{t}p_{j}\log v=\log v\sum_{j=1}^{t}p_{j}=\log v.$ (IV.F.36)
Hence $a$ can just take the value of $\log v$, where $v$ is a positive integer.
As for “if” part of the theorem, it can be proved by Lemma 2 and the fact that $a=\log{v}$ on the ray $\hat{U}^{1}_{3,5}$ and whole ray $U_{1,3}^{234}$ are entropic. ∎
@@
For $F=(\hat{U}_{3,5}^{1},U_{2,3}^{123})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a=\log{v_{1}},b=\log{v_{2}}$ for positive integer $v_{1},v_{2}$.
@@
$H(X_{ik})+H(X_{jk})$ $=H(X_{k})+H(X_{i,j,k}),$
$\qquad\qquad\qquad\quad\{i,j,k\}=\{2,3,4\},$
$H(X_{i\cup K})+H(X_{j\cup K})$ $=H(X_{K})+H(X_{ij\cup K}),$
$\qquad\qquad\qquad\quad K=\{1,2\},\{1,3\}.$
For $(x_{i},i\in N_{4})\in\mathcal{X}_{N_{4}}$ with $p(x_{1234})>0$, above information equalities imply that the probability mass function satisfies
$p(x_{1},x_{2},x_{3},x_{4})$ $=p(x_{1},x_{2},x_{3})$ (IV.F.37)
$=p(x_{1},x_{2},x_{4})$ (IV.F.38)
$=p(x_{1},x_{3},x_{4})$ (IV.F.39)
$=p(x_{2},x_{3},x_{4}),$ (IV.F.40)
$p(x_{1},x_{2})$ $=p(x_{1})p(x_{2}),$ (IV.F.41)
$p(x_{1},x_{3})$ $=p(x_{1})p(x_{3}),$ (IV.F.42)
$p(x_{1},x_{4})$ $=p(x_{1})p(x_{4}),$ (IV.F.43)
$p(x_{2},x_{3})$ $=p(x_{2})p(x_{3}),$ (IV.F.44)
$p(x_{2},x_{4})$ $=p(x_{2})p(x_{4}),$ (IV.F.45)
$p(x_{3},x_{4})$ $=p(x_{3})p(x_{4}),$ (IV.F.46)
$p(x_{2},x_{3})p(x_{2},x_{4})$ $=p(x_{2})p(x_{2},x_{3},x_{4}),$ (IV.F.47)
$p(x_{2},x_{3})p(x_{3},x_{4})$ $=p(x_{3})p(x_{2},x_{3},x_{4}),$ (IV.F.48)
$p(x_{2},x_{4})p(x_{3},x_{4})$ $=p(x_{4})p(x_{2},x_{3},x_{4}),$ (IV.F.49)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{2},x_{4})$ $=p(x_{1},x_{2})p(x_{1234}),$ (IV.F.50)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{3},x_{4})$ $=p(x_{1},x_{3})p(x_{1234}).$ (IV.F.51)
By (IV.F.37)-(IV.F.40), (IV.F.50) and (IV.F.51), we have
$p(x_{1},x_{2},x_{3},x_{4})=p(x_{1},x_{2})=p(x_{1},x_{3}).$ (IV.F.52)
@@
$p(x_{1},x_{2},x_{3},x_{4})=p(x_{1})p(x_{2})=p(x_{1})p(x_{3}),$ (IV.F.53)
which implies $p(x_{2})=p(x_{3})$. Since $X_{2}$ and $X_{3}$ are independent by (IV.F.44), with Lemma 1, $X_{2}$ and $X_{3}$ are uniformly distributed on $\mathcal{X}_{2}$ and $\mathcal{X}_{3}$ and
$H(X_{2})=H(X_{3})=\log v,$ (IV.F.54)
where $v=|\mathcal{X}_{2}|=|\mathcal{X}_{3}|$. Note that $X_{2}$ is independent of $X_{3}$ and $X_{4}$, replacing $p(x_{2},x_{3})$ and $p(x_{2},x_{4})$ by $p(x_{2})p(x_{3})$ and $p(x_{2})p(x_{4})$ in (IV.F.47), respectively, we have
$p(x_{2},x_{3},x_{4})=p(x_{2})p(x_{3})p(x_{4}).$ (IV.F.55)
Since $p(x_{1},x_{2},x_{3},x_{4})=p(x_{2},x_{3},x_{4})$ by (IV.F.40), together with (IV.F.53) and (IV.F.55),
$p(x_{1})=p(x_{3})p(x_{4})=p(x_{2})p(x_{4}).$ (IV.F.56)
Note that $p(x_{2})=p(x_{3})=\frac{1}{v}$, we obtain
$p(x_{1})=\dfrac{1}{v}p(x_{4}).$ (IV.F.57)
@@
$p(x_{1})=\dfrac{1}{v}p(x_{4}^{\prime}).$ (IV.F.59)
Equating (IV.F.57) and (IV.F.59),
$p(x_{4})=p(x_{4}^{\prime}),$ (IV.F.60)
which implies that $X_{4}$ is uniformly distributed on $\mathcal{X}_{4}$. Similarly, $X_{1}$ is uniformly distributed on $\mathcal{X}_{1}$. Assume $|X_{4}|=v_{1}$, we conclude that
$H(X_{1})=\log vv_{1},H(X_{4})=\log v_{1}.$ (IV.F.61)
@@
$p(x_{1},x_{4})=p(x_{1})p(x_{4})=\dfrac{1}{vv_{1}}\times\dfrac{1}{v_{1}}=\dfrac{1}{vv^{2}_{1}}.$ (IV.F.62)
@@
$p(x_{1},x_{4})$ $=\sum_{x_{2},x_{3}}p(x_{1234})=\sum_{x_{2},x_{3}}p(x_{2},x_{3},x_{4})$ (IV.F.63)
$=\sum_{x_{2},x_{3}}p(x_{2})p(x_{3})p(x_{4})=\dfrac{k_{2}}{k^{2}k_{1}},$ (IV.F.64)
where $v_{2}=\mid\{(x_{2},x_{3})|\quad p(x_{1},x_{2},x_{3},x_{4})>0\}\mid$, we conclude that $v=v_{1}v_{2}$. As $\mathbf{h}\in F$ and $(X_{i},i\in N_{4})$ is its characterizing random vector, we have
$H(X_{1})$ $=2a+b,$ (IV.F.65)
$H(X_{2})$ $=H(X_{3})=a+b,$ (IV.F.66)
$H(X_{4})$ $=a.$ (IV.F.67)
Together with (IV.F.54) and (IV.F.61), we obtain
$a=\log v_{1},b=\log\frac{v}{v_{1}}=\log v_{2},$ (IV.F.68)
which forms an outer bound on the entropy region on $F$, that is, $a$ and $b$ must take the value of $\log v_{1}$ and $\log v_{2}$, respectively, where $v_{1}$ and $v_{2}$ are positive integers.
The “if” part of the theorem can be immediately proved using Lemma 2, given the fact that $a=\log{v_{1}}$ in $\hat{U}^{1}_{3,5}$, and $b=\log{v_{2}}$ in $U_{2,3}^{123}$, both of which are entropic. ∎
@@
For $F=(\hat{U}_{3,5}^{1},\mathcal{W}_{2}^{23})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if a+ba+b= $\log{v}$ for integer $v>0$.
@@
$H(X_{i\cup K})+H(X_{j\cup K})$ $=H(X_{K})+H(X_{ij\cup K}),\mid K\mid=2,$
$\hskip 108.12054pt\{1\}\subseteq K.$
For $(x_{i},i\in N_{4})\in\mathcal{X}_{N_{4}}$ with $p(x_{1234})>0$, above information equalities imply that the probability mass function satisfies
$p(x_{1},x_{2},x_{3},x_{4})$ $=p(x_{1},x_{2},x_{3})$ (IV.F.69)
$=p(x_{1},x_{2},x_{4})$ (IV.F.70)
$=p(x_{1},x_{3},x_{4})$ (IV.F.71)
$=p(x_{2},x_{3},x_{4}),$ (IV.F.72)
$p(x_{1},x_{2})$ $=p(x_{1})p(x_{2}),$ (IV.F.73)
$p(x_{1},x_{3})$ $=p(x_{1})p(x_{3}),$ (IV.F.74)
$p(x_{1},x_{4})$ $=p(x_{1})p(x_{4}),$ (IV.F.75)
$p(x_{2},x_{4})$ $=p(x_{2})p(x_{4}),$ (IV.F.76)
$p(x_{3},x_{4})$ $=p(x_{3})p(x_{4}),$ (IV.F.77)
$p(x_{2},x_{3})p(x_{2},x_{4})$ $=p(x_{2})p(x_{2},x_{3},x_{4}),$ (IV.F.78)
$p(x_{2},x_{3})p(x_{3},x_{4})$ $=p(x_{3})p(x_{2},x_{3},x_{4}),$ (IV.F.79)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{2},x_{4})$ $=p(x_{1},x_{2})p(x_{1234}),$ (IV.F.80)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{3},x_{4})$ $=p(x_{1},x_{3})p(x_{1234}),$ (IV.F.81)
$p(x_{1},x_{2},x_{4})p(x_{1},x_{3},x_{4})$ $=p(x_{1},x_{4})p(x_{1234}).$ (IV.F.82)
By (IV.F.69)-(IV.F.72) and (IV.F.80)-(IV.F.82), we have
$p(x_{1234})=p(x_{1},x_{2})=p(x_{1},x_{3})=p(x_{1},x_{4}).$ (IV.F.83)
Since $X_{1}$ is independent of $X_{2},X_{3}$ and $X_{4}$ by (IV.F.73)-(IV.F.75), together with(IV.F.83)
$p(x_{1})p(x_{2})=p(x_{1})p(x_{3})=p(x_{1})p(x_{4}),$ (IV.F.84)
which implies $p(x_{2})=p(x_{3})=p(x_{4})$. Note that $X_{4}$ is independent of $X_{2}$ and $X_{3}$ by (IV.F.76) and (IV.F.77), by Lemma 1, $X_{i}$ is uniformly distributed on $\mathcal{X}_{i}$ and $H(X_{i})=\log v$ for $i=2,3,4$, where $v=|\mathcal{X}_{2}|=|\mathcal{X}_{3}|=|\mathcal{X}_{4}|$. As $\mathbf{h}\in F$ and $(X_{i},i\in N_{4})$ is its characterizing random vector, we have
$H(X_{2})=H(X_{3})=H(X_{4})=a+b.$ (IV.F.85)
Hence $a+b$ must take the value of $\log v$.
As for the “if” part of the theorem, we show that all the polymatroids $(a,b)$ satisfying $a+b=\log v$ are entropic. Let $p_{1},p_{2},\dots,p_{k}$ be positive value satisfying $H(p_{1},p_{2},\dots,p_{k})=a$. Let $X_{1}$ be distributed on $\mathbb{I}_{v^{2}}$ such that
$p_{X_{1}}(i)=\dfrac{p_{\lceil\frac{i+1}{k}\rceil}}{v},\text{ for }i=0,1,...,v^{2}-1.$ (IV.F.86)
Let $X_{2}$ be independent of $X_{1}$ and uniformly distributed on $\mathbb{I}_{v}$. Let
$X_{3}$ $=(\lfloor\frac{X_{1}}{v}\rfloor+X_{2})\mod v,$ (IV.F.87)
$X_{4}$ $=(X_{1}\mod v+X_{2})\mod v.$ (IV.F.88)